OpenStudy (anonymous):

Over 2 years, how much does $2000 in a savings account with an APR of 4.2% compounded semiannually earn in interest than the same amount in a savings account with an APR of 3.6% compounded monthly? A. $6.07 B. $12.15 C. $48.58 D. $24.29

OpenStudy (anonymous):

A = P(1+r/n)^(nt) Use this formula

OpenStudy (anonymous):

P = $2000 r = 0.042 t = 2 A = ? n = 2

OpenStudy (wolf1728):

4.2% Interest Total = 2,000 * (1+ (.042/2)^2*2 Total = 2,000 * 1.0866832385 Total = 2,173.37 3.6% Interest compounded monthly Total = 2,000 * (1+ (.036/12))^24 Total = 2,000 * 1.0745395191 Total = 2,149.08 The difference between the two amounts is 24.29 so the answer is D. Here's a calculator that converts annual interest to semi-annual, quarterly, monthly, daily or continuous interest. http://1728.org/intrate.htm

OpenStudy (anonymous):

Thanks!